# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

### v1

@@ 1 Introduction
In this paper, we study contextual MNL bandit problem [5, 4, 24, 9, 21, 22, 15, 3], where the features of items and possibly contextual information about a user at each round are available. Despite many recent advances, [9, 21, 22, 15, 3], however, no previous studies have proven the minimax optimality of contextual MNL bandits. Chen et al. [9] proposed a regret lower bound of $\Omega(d\sqrt{T}/K)$, where $d$ is the number of features, $T$ is the total number of rounds, assuming the uniform rewards, i.e., $r_{ti}=1$ for all items $i$ and all rounds $t$. Furthermore, Chen and Wang [8] established a regret lower bound of $\Omega(\sqrt{NT})$ in the non-contextual setting (hence, dependence on $N$ appears instead of $d$), which is tighter in terms of $K$. It is important to note the difference in the assumptions for the utility of the outside option $v_{0}$. Chen and Wang [8] assume for the utility for the outside option to be $v_{0}=K$, whereas Chen et al. [9] assume $v_{0}=1$, Therefore, it remains an open question whether and how the value of $v_{0}$ affects both lower and upper bounds of regret.
Regarding upper bounds, Chen et al. [9] proposed an exponential running time algorithm that achieves a regret of $\tilde{\mathcal{O}}(d\sqrt{T})$ in the setting with stochastic contexts and the non-uniform rewards. Under the same setting, Oh and Iyengar [22] introduced a polynomial-time algorithm that attains a regret of $\tilde{\mathcal{O}}(d\sqrt{T}/\kappa)$, where $\kappa=\mathcal{O}(1/K^{2})$ is a problem-dependent constant. Recently, Goyal and Perivier [15] improved the dependency on $\kappa$ in the more challenging adversarial context setting, achieving a regret of $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T})$, where $\kappa^{\prime}=\mathcal{O}(1/K)$. However, their approach focuses solely on the setting with uniform rewards, which is a special case of non-uniform rewards, and currently, there is no tractable method to implement the algorithm.
As summarized in Table 1, there has been a gap between the upper and lower bounds in existing works of contextual MNL bandits. No previous studies have confirmed whether lower or upper bounds are tight, obscuring what the optimal regret should be. This ambiguity is further exacerbated because many studies introduce their methods under varying conditions such as different reward structures and values of $v_{0}$, without explicitly explaining how these factors impact regret. Additionally, there is currently no computationally efficient algorithm whose regret does not scale with $K$, especially in adversarial contexts. Intuitively, increasing $K$ provides more information at least in the uniform reward setting, potentially leading to the more statistically efficient learning process. However, no previous results have reflected such intuition. Hence, the following natural questions arise:
@@
Table 1: Comparison of lower and upper regret bounds in related works on MNL bandits with $T$ rounds, $N$ items, the maximum size of assortments $K$, $d$-dimensional feature vectors, and problem-dependent constants $\kappa=\mathcal{O}(1/K^{2})$ and $\kappa^{\prime}=\mathcal{O}(1/K)$. $\tilde{\mathcal{O}}$ represents big-$\mathcal{O}$ notation up to logarithmic factors. For the computational cost (abbreviated as “Comput.”), we consider only the dependence on the time step $t$. “Intractable” means a non-polynomial runtime. The notation “-” denotes N/A, not applicable. The starred () papers considered the non-contextual setting, which we adapt to the contextual setting by simply setting $d=N$.
@@
Agrawal et al. [5] $\Omega(\sqrt{dT/K})$- Uniform $\Theta(K)$-
Chen and Wang [8] $\Omega(\sqrt{dT})$- Uniform $\Theta(K)$-
@@
Goyal and Perivier [15] $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T})$ Adversarial Uniform $\Theta(1)$ Intractable
@@
  • Under uniform rewards, we establish a regret lower bound of $\Omega(\sqrt{v_{0}K}/(v_{0}+K)d\sqrt{T})$ which is the tightest known lower bound in contextual MNL bandit. We propose, for the first time, a computationally efficient and provably optimal (up to logarithmic factors) algorithm, OFU-MNL+, achieving a matching upper bound of $\mathcal{O}(\sqrt{v_{0}K}/(v_{0}+K)d\sqrt{T})$ up to logarithmic factors while requiring only a constant computation cost per round. The results indicate that the regret improves as the assortment size $K$ increases unless $v_{0}=\Theta(K)$. To the best of our knowledge, this is the first study to demonstrate the dependence of regret on the utility for the outside option $v_{0}$ and to highlight the advantages of a larger assortment size $K$ which aligns with intuition.
  • Under non-uniform rewards, with setting $v_{0}=\Theta(1)$ following the convention in contextual MNL bandits [21, 22, 15, 3, 29], we establish a regret lower bound of $\Omega(d\sqrt{T})$. To the best of our knowledge, this is the first and tightest lower bound established under non-uniform rewards. Moreover, OFU-MNL+ also achieves a matching upper bound (up to logarithmic factors) of $\tilde{\mathcal{O}}(d\sqrt{T})$ in this setting.
@@ 2 Related Work
Lower bounds of MNL bandits. In contextual MNL bandits, to the best of our knowledge, only Chen et al. [9] have proven a lower bound of $\Omega(d\sqrt{T}/K)$ with the utility for the outside option set at $v_{0}=\Theta(1)$. However, in the non-contextual setting, there exist improved lower bounds in terms of $K$.  Agrawal et al. [5] have shown a lower bound of $\Omega(\sqrt{NT/K})$, and Chen and Wang [8] have established a lower bound of $\Omega(\sqrt{NT})$. By simply setting $d=N$, one can derive equivalent lower bounds for the contextual setting, specifically $\Omega(\sqrt{dT/K})$ and $\Omega(\sqrt{dT})$, respectively. However, Agrawal et al. [5] and Chen and Wang [8] assumed that $v_{0}$ is proportional to $K$, i.e., $v_{0}=\Theta(K)$ when proving their lower bounds, which is a different setting from that used in Chen et al. [9], where $v_{0}=\Theta(1)$. Moreover, to the best of our knowledge, all existing works Chen et al. [9], Agrawal et al. [5], Chen and Wang [8] have established the lower bounds under uniform rewards. Consequently, it remains unclear what the optimal regret is, depending on the value of $v_{0}$ and the reward structure.
@@
Upper bounds of contextual MNL bandits. Ou et al. [24] formulated a linear utility model and achieved $\tilde{\mathcal{O}}(dK\sqrt{T})$ regret; however, they assumed that utilities are fixed over time. Chen et al. [9] considered the contextual MNL bandit with changing and stochastic contexts, establishing a regret of $\tilde{\mathcal{O}}(d\sqrt{T}+d^{2}K^{2})$. However, they encountered computational issues due to the need to enumerate all possible ($N$ choose $K$) assortments. To address this, Oh and Iyengar [22] proposed a polynomial-time assortment optimization algorithm, which maintains the confidence bounds in the parameter space and then calculates the upper confidence bounds of utility for each items, achieving a regret of $\tilde{\mathcal{O}}(d\sqrt{T}/\kappa)$, where $\kappa=\mathcal{O}(1/K^{2})$ is a problem-dependent constant. Goyal and Perivier [15] considered the adversarial context and uniform reward setting and improved the dependency on $\kappa$ to $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T}+d^{2}K^{4}/\kappa)$, where $\kappa^{\prime}=\mathcal{O}(1/K)$. However, their algorithm is intractable. Recently, Zhang and Sugiyama [29] utilized an online parameter update to construct a constant time algorithm. However, they consider a multiple-parameter choice model in which the learner estimates $K$ parameters and shares the contextual information $x_{t}$ across the items in the assortment. This model differs from ours; we use a single-parameter choice model with varying the context for each item in the assortment. Additionally, they make a stronger assumption regarding the reward than we do (see Assumption 1). Moreover, while they fix the assortment size at $K$, we allow it to be smaller than or equal to $K$. To the best of our knowledge, all existing methods fail to show that the regret upper bound can improve as the assortment size $K$ increases.
3 Existing Gap between Upper and Lower Bounds in MNL Bandits
The primary objective of this paper is to bridge the existing gap between the upper and lower bounds and to establish minimax regrets in contextual MNL bandits. To explore the optimality of regret, we analyze how it depends on the utility of the outside option $v_{0}$, the maximum assortment size $K$, and the structure of rewards.
[…]
@@ 4 Problem Setting
Notations. For a positive integer, $n$, we denote $[n]:=\{1,2,\ldots,n\}$. For a real-valued matrix $A$, we denotes $\|A\|_{2}:=\sup_{x:\|x\|_{2}=1}\|Ax\|_{2}$ as the maximum singular value of $A$. For two symmetric matrices, $V$ and $W$ of the same dimensions, $V\succeq W$ means that $V-W$ is positive semi-definite. Finally, we define $\mathcal{S}$ to be the set of candidate assortment with size constraint at most $K$, i.e., $\mathcal{S}=\{S\subset[N]:|S|\leq K\}$. While, for simplicity, we consider both $\mathcal{S}$ and the set of items $[N]$ to be stationary in this paper, it is important to note that both $\mathcal{S}$ and $[N]$ can vary over time.
Contextual MNL bandit. We consider a sequential assortment selection problem which is defined as follows. At each round $t$, the agent observes feature vectors $x_{ti}\in\mathbb{R}^{d}$ for every item $i\in[N]$. Based on this contextual information, the agent presents an assortment $S_{t}=\{i_{1},\dots,i_{l}\}\in\mathcal{S}$, where $l\leq K$, and then observes the user purchase decision $c_{t}\in S_{t}\cup\{0\}$, where $\{0\}$ represents the “outside option” which indicates that the user did not select any of the items in $S_{t}$. The distribution of these selections follows a multinomial logit (MNL) choice model [18], where the probability of choosing any item $i_{k}\in S_{t}$ (or the outside option) is defined as:
@@
where $v_{0}$ is a known utility for the outside option and $\mathbf{w}^{\star}\in\mathbb{R}^{d}$ is an unknown parameter. This model ensures that the loss function (to be specified later) has the constant parameter self-concordant-like property (Proposition B.1).
@@
In the existing literature on MNL bandits, it is commonly assumed that $v_{0}=\Theta(1)$ [21, 22, 15, 3, 29]. On the other hand, Chen and Wang [8], Agrawal et al. [5] assume that $v_{0}=\Theta(K)$ 111 Chen and Wang [8] indeed set $v_{0}=1$ and $v_{1},\dots,v_{N}=\Theta(1/K)$. However, this is equivalent to setting $v_{0}=\Theta(K)$ and $v_{1},\dots,v_{N}=\Theta(1)$. to induce a tighter lower bound in terms of $K$. Later, we will explore how these differing assumptions create fundamentally different problems, leading to different regret lower bounds (Subsection 5.1).
@@
and define $S_{t}^{\star}$ as the offline optimal assortment at time $t$ when $\mathbf{w}^{\star}$ is known a prior, i.e., $S_{t}^{\star}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}\sum_{i\in S}R_{t}(S,\mathbf{w}^{\star})$. Our objective is to minimize the cumulative regret over the $T$ periods:
@@
When $K=1$, $r_{t1}=1$, and $v_{0}=1$, MNL bandit recovers the binary logistic bandit with $R_{t}(S=\{x\},\mathbf{w}^{\star})=\sigma\left(x^{\top}\mathbf{w}^{\star}\right)=1/(1+\exp(-x^{\top}\mathbf{w}^{\star}))$, where $\sigma(\cdot)$ is the sigmoid function.
Consistent with previous works on MNL bandit [22, 15, 3, 29], we make the following assumptions:
@@
Assumption 2 is common in contextual MNL bandit [9, 22, 15, 29]. Note that $1/\kappa$ depends on the maximum size of the assortment $K$, i.e., $1/\kappa=\mathcal{O}(K^{2})$. Since one of the primary goals of this paper is to show that, as the assortment size $K$ increases, we can achieve an improved (or at least not worsened) regret bound, we design a dynamic assortment policy that enjoys improved dependence on $\kappa$. Note that our algorithm does not need to know $\kappa$ a priori, whereas Oh and Iyengar [21, 22] do.

### v6

@@ 1 Introduction
In this paper, we study the contextual MNL bandit problem [7, 6, 42, 14, 39, 40, 43, 4], where the features of items and possibly contextual information about a user at each round are available. Despite many recent advances, [14, 39, 40, 43, 4], however, no previous studies have proven the minimax optimality of contextual MNL bandits. Chen et al. [14] proposed a regret lower bound of $\Omega(d\sqrt{T}/K)$, where $d$ is the number of features, $T$ is the total number of rounds, and $K$ is the maximum size of assortments, assuming the uniform rewards, i.e., rewards are all same for each of the total $N$ items. Furthermore, Chen and Wang [13] established a regret lower bound of $\Omega(\sqrt{NT})$ in the non-contextual setting (hence, dependence on $N$ appears instead of $d$), which is tighter in terms of $K$. It is important to note the difference in the assumptions for the attraction parameter for the outside option $v_{0}$. Chen and Wang [13] assumed for the attraction parameter for the outside option to be $v_{0}=K$, whereas Chen et al. [14] assumed $v_{0}=1$. Therefore, it remains an open question whether and how the value of $v_{0}$ affects both lower and upper bounds of regret.
Regarding regret upper bounds, Chen et al. [14] proposed an exponential runtime algorithm that achieves a regret of $\tilde{\mathcal{O}}(d\sqrt{T})$ in the setting with stochastic contexts and the non-uniform rewards. Under the same setting, Oh and Iyengar [40] and Oh and Iyengar [39] introduced polynomial-time algorithms that attain regrets of $\tilde{\mathcal{O}}(d\sqrt{T}/\kappa)$ and $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T}/\kappa)$ respectively, where $1/\kappa=\mathcal{O}(K^{2})$ is a problem-dependent constant. Recently, Perivier and Goyal [43] improved the dependency on $\kappa$ in the adversarial context setting, achieving a regret of $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T})$, where $\kappa^{\prime}=\mathcal{O}(1/K)$. However, their approach focuses solely on the setting with uniform rewards, which is a special case of non-uniform rewards, and currently, there is no tractable method to implement the algorithm.
As summarized in Table 1, there has been a gap between the upper and lower bounds in the existing works of contextual MNL bandits. No previous studies have confirmed whether lower or upper bounds are tight, obscuring what the optimal regret should be. This ambiguity is further exacerbated because many studies introduce their methods under varying conditions such as different reward structures and values of $v_{0}$, without explicitly explaining how these factors impact regret. Additionally, there is currently no computationally efficient algorithm whose regret does not scale with $1/\kappa=\mathcal{O}(K^{2})$ or directly with $K$. Intuitively, increasing $K$ provides more information at least in the uniform reward setting, potentially leading to a more statistically efficient learning process. However, no previous results have reflected such intuition. Hence, the following research questions arise:
@@
Table 1: Comparisons of lower and upper regret bounds in related works on MNL bandits with $T$ rounds, $N$ items, the maximum size of assortments $K$, $d$-dimensional feature vectors, and problem-dependent constants $1/\kappa=\mathcal{O}(K^{2})$ and $\kappa^{\prime}=\mathcal{O}(1/K)$. $\tilde{\mathcal{O}}$ represents big-$\mathcal{O}$ notation up to logarithmic factors. For the computational cost (abbreviated as “Comput.”), we consider only the dependence on the number of rounds $t$. “Intractable” means a non-polynomial runtime. The notation “--” denotes not applicable. The starred () papers only consider the non-contextual setting.
@@
Agrawal et al. [7] $\Omega(\sqrt{NT/K})$- Uniform $\Theta(K)$-
Chen and Wang [13] $\Omega(\sqrt{NT})$- Uniform $\Theta(K)$-
@@
Oh and Iyengar [39] $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T}/\kappa)$ Adversarial Non-uniform $\Theta(1)$ $\mathcal{O}(t)$
Perivier and Goyal [43] $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T})$ Adversarial Uniform $\Theta(1)$ Intractable
@@
  • Under uniform rewards, we establish a regret lower bound of $\Omega(\sqrt{v_{0}K}/(v_{0}+K)d\sqrt{T})$ (Theorem 1), which is the tightest known lower bound in contextual MNL bandits. We propose, for the first time, a computationally efficient and provably optimal algorithm, OFU-MNL+, achieving a matching upper bound of $\mathcal{O}(\sqrt{v_{0}K}/(v_{0}+K)d\sqrt{T})$ (Theorem 2) up to logarithmic factors, while requiring only a constant computation cost per round. The results indicate that the regret improves as the assortment size $K$ increases, unless $v_{0}=\Theta(K)$. To the best of our knowledge, this is the first study to demonstrate the dependence of regret on the attraction parameter for the outside option $v_{0}$ and to highlight the advantages of a larger assortment size $K$ which aligns with intuition. That is, this is the first work to show that a regret upper bound (in either contextual or non-contextual setting) decreases as $K$ increases. Additionally, we provide instance-dependent minimax regret bounds (Proposition 1 and 2), up to logarithmic factors.
  • Under non-uniform rewards, with setting $v_{0}=\Theta(1)$ following the convention in contextual MNL bandits [5, 16, 42, 7, 39, 40, 8, 43, 4, 52, 32], we establish a regret lower bound of $\Omega(d\sqrt{T})$ (Theorem 3). To the best of our knowledge, this is the first and tightest lower bound established under non-uniform rewards. Moreover, OFU-MNL+ also achieves a matching upper bound (up to logarithmic factors) of $\tilde{\mathcal{O}}(d\sqrt{T})$ (Theorem 4) in this setting.
@@
Overall, our paper addresses the long-standing open problem of closing the gap between upper and lower bounds for contextual MNL bandits. Our proposed algorithm is the first to achieve both provably optimality (up to logarithmic factors) and practicality with improved computation.
@@ 2 Related Work
Lower bounds of MNL bandits. In contextual MNL bandits, to the best of our knowledge, only Chen et al. [14] proved a lower bound of $\Omega(d\sqrt{T}/K)$ with the attraction parameter for the outside option set at $v_{0}=1$. However, in the non-contextual setting, there exist improved lower bounds in terms of $K$.  Agrawal et al. [7] demonstrated a lower bound of $\Omega(\sqrt{NT/K})$, and Chen and Wang [13] established a lower bound of $\Omega(\sqrt{NT})$. By setting $d=N$, one can derive equivalent lower bounds for the contextual setting, specifically $\Omega(\sqrt{dT/K})$ and $\Omega(\sqrt{dT})$, respectively. However, Agrawal et al. [7] and Chen and Wang [13] assumed $v_{0}=K$ when establishing their lower bounds, which differs from the setting used by Chen et al. [14], where $v_{0}=1$. Moreover, to the best of our knowledge, all existing works Chen et al. [14], Agrawal et al. [7], Chen and Wang [13] have established the lower bounds under uniform rewards. Consequently, it remains unclear what the optimal regret is, depending on the value of $v_{0}$ and the reward structure.
@@
Upper bounds of contextual MNL bandits. Ou et al. [42] formulated a linear utility model and achieved $\tilde{\mathcal{O}}(dK\sqrt{T})$ regret; however, they assumed that utilities are fixed over time. Chen et al. [14] considered contextual MNL bandits with changing and stochastic contexts, establishing a regret of $\tilde{\mathcal{O}}(d\sqrt{T}+d^{2}K^{2})$. However, they encountered computational issues due to the need to enumerate all possible ($N$ choose $K$) assortments. To address this, Oh and Iyengar [40] proposed a polynomial-time assortment optimization algorithm, which maintains the confidence bounds in the parameter space and then calculates the upper confidence bounds of attraction parameter for each item, achieving a regret of $\tilde{\mathcal{O}}(d\sqrt{T}/\kappa)$, where $1/\kappa=\mathcal{O}(K^{2})$ is a problem-dependent constant. Perivier and Goyal [43] considered the adversarial context and uniform reward setting and improved the dependency on $\kappa$ to $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T}+d^{2}K^{4}/\kappa)$, where $\kappa^{\prime}=\mathcal{O}(1/K)$. However, their algorithm is intractable. Agrawal et al. [4] considered a uniform rewards setting (with $v_{0}=1$) and achieved a regret of $\tilde{\mathcal{O}}(d\sqrt{T})$. However, due to significant technical errors in their paper (refer Appendix L), we do not include a comparison with their results in this work.
Recently, Zhang and Sugiyama [52] utilized an online parameter update to construct a constant time algorithm. However, they consider a multiple-parameter choice model in which the learner estimates $K$ parameters and shares the contextual information $x_{t}$ across the items in the assortment. This model differs from ours; we use a single-parameter choice model with varying the context for each item in the assortment. Additionally, they make a stronger assumption regarding the reward than we do (see Assumption 1). Moreover, while they fix the assortment size at $K$, we allow it to be smaller than or equal to $K$. On the other hand, Zhang and Luo [51] considered a general function approximation, achieving a regret bound of $\tilde{\mathcal{O}}(K^{2.5}\sqrt{dNT})$. However, this bound scales with $K$ and $N$, and the proposed algorithm is not tractable. To the best of our knowledge, all existing methods fail to show that the regret upper bound can improve as the assortment size $K$ increases.
@@ 3 Problem Setting
Notations. For a positive integer, $n$, we denote $[n]:=\{1,2,\ldots,n\}$. For a real-valued matrix $A$, we denote $\|A\|_{2}:=\sup_{x:\|x\|_{2}=1}\|Ax\|_{2}$ as the maximum singular value of $A$. For two symmetric matrices, $V$ and $W$ of the same dimensions, $V\succeq W$ means that $V-W$ is positive semi-definite. Finally, we define $\mathcal{S}$ to be the set of candidate assortment with size constraint at most $K$, i.e., $\mathcal{S}=\{S\subset[N]:|S|\leq K\}$. While, for simplicity, we consider both $\mathcal{S}$ and the set of items $[N]$ to be stationary in this paper, it is important to note that both $\mathcal{S}$ and $[N]$ can vary over time.
Contextual MNL bandits. We consider a sequential assortment selection problem which is defined as follows. At each round $t$, the agent observes feature vectors $x_{ti}\in\mathbb{R}^{d}$ for every item $i\in[N]$. Based on this contextual information, the agent presents an assortment $S_{t}=\{i_{1},\dots,i_{l}\}\in\mathcal{S}$, where $l\leq K$, and then observes the user purchase decision $c_{t}\in S_{t}\cup\{0\}$, where $\{0\}$ represents the “outside option” which indicates that the user did not select any of the items in $S_{t}$. The distribution of these selections follows a multinomial logit (MNL) choice model [36], where the probability of choosing any item $i_{k}\in S_{t}$ (or the outside option) is defined as:
@@
where $v_{0}$ is a known attraction parameter for the outside option and $\mathbf{w}^{\star}\in\mathbb{R}^{d}$ is an unknown parameter.
@@
In the existing literature on MNL bandits, it is commonly assumed that $v_{0}=1$ [39, 40, 43, 4, 52]. On the other hand, Chen and Wang [13], Agrawal et al. [7] assume that $v_{0}=K$ 111 Chen and Wang [13] indeed set $v_{0}=1$ and $v_{1},\dots,v_{N}=\Theta(1/K)$. However, this is equivalent to the setting with $v_{0}=K$ and $v_{1},\dots,v_{N}=\Theta(1)$. to induce a tighter lower bound in terms of $K$. Later, we will explore how these differing assumptions create fundamentally different problems, leading to different regret lower bounds (Subsection 5.1).
@@
and define $S_{t}^{\star}$ as the offline optimal assortment at time $t$ when $\mathbf{w}^{\star}$ is known a prior, i.e., $S_{t}^{\star}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}R_{t}(S,\mathbf{w}^{\star})$. Our objective is to minimize the cumulative regret over the $T$ periods:
@@
When $K=1$, $r_{t1}=1$, and $v_{0}=1$, the MNL bandit recovers the binary logistic bandit with $R_{t}(S=\{x\},\mathbf{w}^{\star})=\sigma\left(x^{\top}\mathbf{w}^{\star}\right)=1/(1+\exp(-x^{\top}\mathbf{w}^{\star}))$, where $\sigma(\cdot)$ is the sigmoid function.
Consistent with previous works on MNL bandits [40, 43, 4, 52], we make the following assumptions:
@@
Assumption 2 is common in contextual MNL bandits [14, 40, 43, 52]. Note that $1/\kappa$ depends on the maximum size of the assortment $K$, i.e., $1/\kappa=\mathcal{O}(K^{2})$. One of the primary goals of this paper is to show that as the assortment size $K$ increases, we can achieve an improved (or at least not worsened) regret bound. To this end, we design a dynamic assortment policy that enjoys improved dependence on $\kappa$. Note that our algorithm does not need to know $\kappa$ a priori, whereas Oh and Iyengar [39, 40] do.
4 Existing Gap between Upper and Lower Bounds in MNL Bandits
The primary objective of this paper is to establish minimax regrets in contextual MNL bandits. To explore the optimality of regret, we analyze how it depends on the attraction parameter for the outside option $v_{0}$, the maximum assortment size $K$, and the structure of rewards.
[…]